Disease Prevalence 50%

Prevalence Disease 10%

Prevalence Disease 1%

Yes

No

Yes

No

Yes

No

Test YES

TP   800 A

FP 100 B

TP   160 A

FP 180 B

TP 16   A

FP    198 B

Test NO

FN 200 C

TN 900 D

FN     40 C

TN 1,620 D

   FN    4   C

TN 1,782 D

Totals

1,000

1,000

200

1,800

20

1,980

NPV=   TN    OR   D    =

          FN+TN      C + D

X 100- ________%

NPV=

NPV=
Answer #1

TP - true positive means both test and disease is positive

TN - true negative means both test and disease is negative

FP - false positive means test is positive ,but disease is absent

FN - false negative means test is negative ,but disease is present

NPV is Negative predictive value . It is equal to true negative divided by sum of false negative and true negative

In disease prevalence 50% ,the TN = 900 ,FN =200

NPV = TN/FN+TN = 900/ 900+200 = 900/1100

= 0.81

In disease prevalence 10% , TN = 1620, FN = 40

NPV = TN/FN+TN = 1620+/1620+40

= 1620/ 1660

=0.97

In disease prevalence 1% ,TN = 1782, FN = 4

NPV = TN/ TN +FN = 1783/1782+4

= 1782/1786

= 0.99